OpenStudy (darkbeautystar):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

OpenStudy (darkbeautystar):

Please help @hartnn

OpenStudy (darkbeautystar):

I need help

OpenStudy (gabebae):

Hello

OpenStudy (darkbeautystar):

Hay

OpenStudy (gabebae):

I'll help you

OpenStudy (darkbeautystar):

Well my question is up top

OpenStudy (gabebae):

A): 0*0=0 0*1 = 0 1*1 = 1 Closed B): 1/1=1 0/1 = 0 Not defined so it's not closed C: 1+1 = 2 Not closed D: 0-1=-1 Not closed Your answer would be A. multiplication

OpenStudy (darkbeautystar):

Ok thanks

OpenStudy (gabebae):

You're welcome

OpenStudy (darkbeautystar):

Here is the next one

OpenStudy (darkbeautystar):

Which equations show that the set of whole numbers is closed under multiplication? Choose all answers that are correct. A. –1 • –1 = 1 B. 0 • –1 = 0 C. 0 • 2 = 0 D. 2 • 1 = 2

OpenStudy (gabebae):

There's more?

OpenStudy (darkbeautystar):

there is 5 questions???

OpenStudy (darkbeautystar):

there are 5 questions

OpenStudy (gabebae):

I would go with A and D on this one

OpenStudy (darkbeautystar):

OK can you tell me why

OpenStudy (gabebae):

hold on I take that back Do you know what whole numbers are?

OpenStudy (darkbeautystar):

Well..... like 1-2-3-4-5

OpenStudy (darkbeautystar):

I am not really sure

OpenStudy (gabebae):

Whole numbers can be positive and negative

OpenStudy (darkbeautystar):

ok

OpenStudy (gabebae):

0 is a whole number as well

OpenStudy (darkbeautystar):

k

OpenStudy (gabebae):

now look at your answer choices

OpenStudy (gabebae):

The answer would be all of them

OpenStudy (darkbeautystar):

because why

OpenStudy (gabebae):

Whole numbers don't have fractions or decimal parts and NO negatives You confused me when you only added negatives sorry

OpenStudy (darkbeautystar):

its ok

OpenStudy (gabebae):

next question?

OpenStudy (darkbeautystar):

Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (darkbeautystar):

are u good in math

OpenStudy (darkbeautystar):

hello welshfella

OpenStudy (gabebae):

If you add any two numbers in a set and you get another number in that set then the set is closed under addition Here's an example 2+2=4 2+4=6

OpenStudy (gabebae):

Your two answers would be B and C

OpenStudy (darkbeautystar):

#4

OpenStudy (darkbeautystar):

Numbers p and q are whole numbers. Which statements are always true? Choose all answers that are correct. A. p + q is a whole number B. p – q is a whole number C. p • q is a whole number D. is a whole number

OpenStudy (gabebae):

What's D?

OpenStudy (darkbeautystar):

positive and negative number

OpenStudy (darkbeautystar):

???

OpenStudy (gabebae):

It got cut off

OpenStudy (darkbeautystar):

Ohhh D. is a whole number

OpenStudy (gabebae):

what is a whole number?

OpenStudy (darkbeautystar):

positive and negative number

OpenStudy (gabebae):

It's A and C

OpenStudy (darkbeautystar):

thanks--last question

OpenStudy (darkbeautystar):

last q and then i will tell you what i got

OpenStudy (gabebae):

okay

OpenStudy (darkbeautystar):

Is the set of multiples of 4 closed under addition? Explain why or provide a counterexample if not. A. Yes, because the sum of any two multiples of 4 is also a multiple of 4. B. Yes, because the sum of any two multiples of 4 is also a multiple of 8. C. No, and a counterexample is –4 + (–8) = –12. D. No, and a counterexample is 26 + 16 = 42.

OpenStudy (gabebae):

A.

OpenStudy (gabebae):

Answer is A

OpenStudy (darkbeautystar):

thanksyou-i will tell you what i got

OpenStudy (gabebae):

Alright

OpenStudy (darkbeautystar):

i got a 65%--not mastered-- :(

OpenStudy (gabebae):

what questions were wrong?

OpenStudy (gabebae):

I could have sworn I knew all of this from algebra

OpenStudy (darkbeautystar):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication <----This one was correct B. division <----this one was correct C. addition D. subtraction

OpenStudy (darkbeautystar):

That was number 1

OpenStudy (gabebae):

Ooh I thought we could only pick one answer

OpenStudy (darkbeautystar):

it says choose all that are correct

OpenStudy (gabebae):

That was the first question you gave me

OpenStudy (gabebae):

and you got another wrong?

OpenStudy (darkbeautystar):

#2. Only C and D were correct Which equations show that the set of whole numbers is closed under multiplication? Choose all answers that are correct. A. –1 • –1 = 1 B. 0 • –1 = 0 C. 0 • 2 = 0 D. 2 • 1 = 2

OpenStudy (gabebae):

meh

OpenStudy (gabebae):

I'm sorry

OpenStudy (gabebae):

I'm more of a geometry person I took algebra last year

OpenStudy (gabebae):

Welshfella should have helped you

OpenStudy (darkbeautystar):

#3. B, C and D were correct Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (gabebae):

Oh

OpenStudy (darkbeautystar):

Thanks anyway

OpenStudy (gabebae):

Gah I feel so dumb now i'm really sorry

OpenStudy (darkbeautystar):

It's ok

OpenStudy (darkbeautystar):

Bye

OpenStudy (gabebae):

gave you a metal